OpenStudy (twizttiez):

A cone has a base radius of 5 centimeters and a height of 18 centimeters. What is the volume of the cone? A. 30π cm3 B. 450π cm3 C. 90π cm3 D. 150π cm3

OpenStudy (twizttiez):

@pooja195

OpenStudy (twizttiez):

I got 471.24

pooja195 (pooja195):

\[V=(3.14)(5^2)\frac{ 18 }{3 }\]

OpenStudy (twizttiez):

Yeah I got 471 again

pooja195 (pooja195):

Oh this doesnt include pi :/

OpenStudy (twizttiez):

Ohhh then what do I do?

OpenStudy (zehanz):

Use the formula for volume: \(V=\frac{1}{3}\cdot A\cdot h\)

OpenStudy (zehanz):

Because the base of the cone is a circle, A (area) is \(\pi \cdot 5^2\). The height h is 18 cm, so put the numbers in the formula to get the right answer...

OpenStudy (twizttiez):

Okay one second

OpenStudy (twizttiez):

My area is 78.5?

OpenStudy (zehanz):

How did you do that?

OpenStudy (twizttiez):

3.14*5^2

OpenStudy (twizttiez):

Right?

OpenStudy (zehanz):

Let's see, we have to calculate the product of: \(\frac{1}{3}\), \(5^2\), \(\pi\) and 18. It is easy to do \(\frac{1}{3}\cdot 18=6\). So now I have \(6\cdot 5^2\cdot \pi=6\cdot 25\cdot \pi=150\pi\). You don't need a calculator, it can be done in the head!

OpenStudy (twizttiez):

Ohhhhhh

OpenStudy (twizttiez):

I kept over thinking it

OpenStudy (twizttiez):

Thanks so much man

OpenStudy (twizttiez):

Could you help me on another one?

OpenStudy (zehanz):

Sorry, have to go. Tomorrow I rise early to go to work...(it's 11.30 PM here).